OpenStudy (anonymous):

M ∠C=75°. Determine the measure of the supplement to ∠C

OpenStudy (amistre64):

complements = 90 while supples = 180

OpenStudy (anonymous):

75+C = 180 C = 180-75 angle C = 105degre